# SATA-Bench: Select All That Apply Benchmark for Multiple Choice Questions

### v2

@@ Appendix B Dataset Filtering
The Biomedicine, Law, and Events datasets were originally multi-label classification tasks, which we adapted into SATA questions by creating distractor (incorrect) choices from the unselected labels. There are two challenges when converting these datasets to SATA format: 1. Many of them have a large label pool with only a few correct answers, which is not reasonable for multiple-choice questions. 2. There can be some popular answers frequently exist in the original data, making certain choices more frequent than others. This is problematic because it may force the model to learn the popular token of choice (e.g. Choice A/a/1) rather than the content of the questions. For example, the law dataset is originally from EUR-Lex data [18] contains 57k legislative documents in English (https://eur-lex.europa.eu) annotated by the Publications Office of EU (https://publications.europa.eu/en) with over 7k concepts from EUROVOC (http://eurovoc.europa.eu/). To address the first challenge, we kept an option-to-answer ratio between 2 and 3, considering the balance between the number of correct answers and incorrect choices. The distractors were sampled randomly from the topic pool. We also shuffled and and re-labeled the choices using unique mapping per question to keep the final answers evenly distributed between each choice token. An example question from each data source is shown in Figure 4.
@@ B.2 Human Validation
Human validation is to ensure that the questions are unambiguous. Using humans to validate the question is inspired by  [49, 33]. For each question in the benchmark, we ask five annotators whether the question contains ambiguous information.
@@
Once it is done, the total cost is tracked (1301.89), with 5 people per label at a cost of 0.012 each. We only select questions that are “Strongly agree” and “Agree" $>0.8$.
B.3 Human Labeling
To ensure that each question has a valid and correct answer, we conducted a comprehensive human evaluation. An initial manual inspection revealed that some questions lacked clearly correct answers. To verify answer correctness, we recruited three experienced annotators to review all 1,650 questions that remained after prior filtering and validation. Annotators were compensated at a rate of at least \$35 per hour. Each question was independently evaluated by at least two annotators.
@@ B.3 Human Labeling
Human Labeling Given original answers and LLMs’ answers, you’ll try to identify correct answer of the following questions. You’re expected/encouraged to use Google, and any internet resources you can find to try and answer the question correctly. Requirements and Expectations 1. You are encouraged to use Google, and any websites you can think of or find that may help you answer the question and understand the concept. However, you are NOT allowed to use AI assistants like chatGPT, Claude, Grok3 Geimini, etc., or ask people for help. All their answers to the question has been provided anoymously under LLM Answers. 2. We ask that you spend at least 5 minutes trying to answer each question before making your selection. If you haven’t settled on an answer choice in that time, we encourage you to spend as long as you need to be confident in your selection. 3. These questions will be hard, and you will likely need to spend a while on each of them to make progress understanding the context. Read relevant resources, take plenty of time, and answer "I don’t know" if you’re pretty sure you have no realistic way of answering confidently. 4. You will also be given the opportunity to give feedback on the question. We’re especially interested in feedback about whether the question was ambiguous, but please feel free to give feedback of any other form! Suggestions and Strategies for Labeling 1. Look up definitions for all of the unfamiliar terms in the question and answer choices. Keep a list of those definitions handy so you can easily refer back to the definitions if you forget the jargon. 2. LLMs’ answer is not always reliable and original answer is not always correct. Please try to solve the question independently before looking at potential answers. 2. Look for primary resources, like research papers and textbooks, as these can often contain clearer explanations than sources like Wikipedia (although Wikipedia can be useful in many cases as well). You are presented with the following: Paragraph: paragraph Question: question Choices: choice Original Answers: original answer LLM Answers: llm answers Answers:
@@ Appendix C Hyperparameters
GPT-4-turbo OpenAI gpt-4o-2024-11-20 11/20/24 OpenAI API
Claude-3.7 Sonnet Think Anthropic anthropic.claude-3-7-sonnet-thinking-20250219-v1:0 02/24/25 AWS Bedrock
Claude-3.7 Sonnet Anthropic anthropic.claude-3-7-sonnet-20250219-v1:0 02/24/25 AWS Bedrock
Claude-3 Sonnet Anthropic anthropic.claude-3-sonnet-20240229-v1:0 02/29/24 AWS Bedrock
@@
Claude-3.5 Haiku Anthropic anthropic.claude-3-5-haiku-20241022-v1:0 10/22/24 AWS Bedrock
Claude-3 Haiku Anthropic anthropic.claude-3-haiku-20240307-v1:0 03/07/24 AWS Bedrock
Claude-3 Opus Anthropic anthropic.claude-3-opus-20240229-v1:0 02/29/24 AWS Bedrock
@@
GPT-4.1 mini OpenAI gpt-4.1-mini-2025-04-14 04/14/25 OpenAI API
Claude-3.5 Sonnet Anthropic anthropic.claude-3-5-sonnet-20240620-v1:0 06/20/24 AWS Bedrock
@@
Phi-3 7B Microsoft microsoft/phi-3-small-128k-instruct 05/21/24 Hugging Face
Phi-4-mini-reasoning Microsoft microsoft/phi-4-mini-reasoning 04/15/25 Hugging Face
@@ Appendix D Compute Resources
We use AWS Bedrock batch inference for large models’ inference such as Claude3 Sonnet, Claude 3.5 Haiku, Claude 3 Haiku, Claude 3 Opus, Claude 3.5 Sonnet, Llama 3.1 405B, Mistral Large V2, Llama 3.2 90B, and Llama 3.1 70B. We use AWS cross-region inference for Claude3.7 Reason, Claude3.7, and Deepseek R1. We use official APIs from the respective providers for models such as OpenAI O3, GPT4.1, Grok3 Reason, GPT4, Geimini2.5 Reason, Gemini 2 Flash, GPT 4.1 mini, and Qwen Plus.
@@ Appendix E Non-expert Human Benchmark
To contextualise LLM results on SATA-Bench, we recruited non-expert annotators on Amazon Mechanical Turk, adapting the instructions from  [43]. All 1604 questions was labelled as follows:
[…]
  • Quality safeguards. Workers were: (i) informed that every item contains at least two correct answers; (ii) forbidden from consulting LLMs or other people, yet allowed to look up unfamiliar terms on Google/Wikipedia; (iii) required to spend $\geq$2 minutes on each question. Submissions that selected any decoy, took $<1$ min, or violated the lookup policy were discarded (7.1 %).
  • Label selection. From the surviving pool, we randomly drew one annotation as the human label; single-choice answers were retained to keep the evaluation comparable to LLMs that sometimes return only one option.
@@
Human 17.9 60.6 54.4 45.0 15.3 0.46 1.46 $-0.6$ 34.1 1.44
Table 8: Aggregate performance of crowd annotators on the SATA-Bench subset.
[…]
@@ Appendix F Metrics Definition
Performance. The standard SATA performance metrics [48] are Exact Match (EM), Jaccard Index (JI), Mean Average Precision (Precision), and Mean Average Recall (Recall). EM captures a model’s ability to select all correct answers without error, indicating its completeness in prediction. JI measures the fraction of predicted labels that match the ground-truth labels.
Selection Bias. Selection bias is a model’s preference for selecting specific option IDs as answers, and is measured by RStd [56] and RSD [11, 42]. We also observe that some models prefer to avoid selecting certain options, which we call unselection bias. We introduce a metric called Selection Probability Divergence (SPD) to measure this type of bias. Appendix G gives more details on this issue and the design of SPD.
Count Bias. We observe that models tend to select fewer number of options compared to the ground truth. We refer to this as Count Bias. To evaluate the severity of this type of bias, we measure the following: (i) Mean difference between the number of selected options minus correct options (CtDif), (ii) Mean absolute difference between the number of selected options minus correct options (CtDifAbs), and (iii) Percentage of cases where the model selects exactly the correct number of options (CtAcc).
@@ F.2 Selection Bias Metrics Definition
    where $k$ is the number of choices, $r_{i}$ is the recall of the $i$-th class, and $\bar{r}$ is the arithmetic mean of $r_{i}$ values. Note that our recalls are calculated at the label level since this is multi-class question [56]
@@
    where $k$ is the number of choices, $s_{i}$ is the accuracy of the $i$-th class, and $\bar{s}$ is the mean accuracy averaged across classes. Please note that our recalls are calculated at the label level since this is multi-class questions [11, 42]
@@ Appendix G Unselection bias metric
SPD has a minimal value of 00 at $q_{i}=p_{i}$ for all $i$, when the prediction aligns with the ground truth. SPD diverges as $q_{i}\rightarrow 0$ while $p_{i}$ is finite for any $i$, when the LLM shows a non-selection bias against a particular label. SPD also diverges as $p_{i}\rightarrow 0$ while $q_{i}$ is finite for any $i$, when the LLM shows a selection bias toward a particular label. In this sense, SPD serves as a metric to measure the disagreement of choice probability between the ground truth and the prediction, reflecting both under-selection and over-selection. (See Appendix G.2 for the mathematical analysis.)
@@ G.1 Behavior of SPD Metric
We repeat the above sampling process $M$ times, and compute the True rate of each option for the ground truth ${p_{i}}$ and the prediction ${q_{i}}$, with $i=1,2,3,4$. We then substitute the numbers into Eq. (3) to calculate SPD. Note that in the current setting, $p_{i}=\mathtt{p},\forall i$, and $q_{1}=\mathtt{q}$, $q_{2,3,4}=\mathtt{p}$.
@@
Figure 8: Relationship between Selection Probability Divergence (SPD) and prediction probability ($\mathtt{q}$) across different ground truth probabilities ($\mathtt{p}$). The curves are averaged over 100 replicates, and the shaded area represents the standard deviation. In each plot, the minimal value of SPD is 00 at $\mathtt{q}=\mathtt{p}$, when the prediction aligns with the ground truth.
@@ G.2 Sensitivity of SPD to label probability ratio
We analyze the behavior of SPD as the relationship between ${p_{i}}$ and ${q_{i}}$ changes. We first define the ratio of the two probabilities as $r_{i}\equiv q_{i}/p_{i}$, $i=1,2,\ldots,k$, and rewrite the SPD definition Eq. (3) as
@@
As the misalignment between the ground truth and the prediction grows, either with $r_{i}\rightarrow 0$ or $r_{i}\rightarrow+\infty$, SPD diverges according to Eq. (4). Therefore, a large value of SPD reflects the disagreement of the choice probability between the ground truth and the prediction.
@@ H.2.1 Problems for current MCQ prompts
Existing benchmarks [25] use the following prompts for MCQ questions and then use exact match to get the correct option.
@@
We can observe that parsing the above prompts is much harder because we do not know when they will output the option ID plus answers, and we cannot blindly extract all matched option IDs. Following [25] page 38, we used the JSON format to extract answers. Closed-source LLMs have shown good instruction following capability for JSON format output [57]. Also, python has many packages that can do fuzzy JSON matching, which can be used to handle edge cases. We extract as many JSON files as possible from each output and combine the extracted answers together as the final answer to the question.
[…]

### v3

@@ Appendix B Dataset Filtering
The Biomedicine, Law, and Events datasets were originally multi-label classification tasks, which we adapted into SATA questions by creating distractor (incorrect) choices from the unselected labels. There are two challenges when converting these datasets to SATA format: 1. Many of them have a large label pool with only a few correct answers, which is not reasonable for multiple-choice questions. 2. There can be some popular answers frequently exist in the original data, making certain choices more frequent than others. This is problematic because it may force the model to learn the popular token of choice (e.g. Choice A/a/1) rather than the content of the questions. For example, the law dataset is originally from EUR-Lex data EUR-Lex (2018) contains 57k legislative documents in English (https://eur-lex.europa.eu) annotated by the Publications Office of EU (https://publications.europa.eu/en) with over 7k concepts from EUROVOC (http://eurovoc.europa.eu/). To address the first challenge, we kept an option-to-answer ratio between 2 and 3, considering the balance between the number of correct answers and incorrect choices. The distractors were sampled randomly from the topic pool. We also shuffled and and re-labeled the choices using unique mapping per question to keep the final answers evenly distributed between each choice token. An example question from each data source is shown in Figure 4.
@@
Table 5: Statistics of the SATA-Bench evaluation dataset (by data source). We report the following metrics: n: number of instances, LC: label cardinality, m: mean number of correct answers, me: median number of correct answers, min: minimum number of correct answers, max: maximum number of correct answers, r: ratio of the number of choices to the median number of correct answers (LC/me), w: mean word count, FRE: Flesch Reading Ease score, FGL: Flesch-Kincaid Grade Level score, ARI: Automated Readability Index, DCR: Dale-Chall Readability score, GFI: Gunning Fog Index, Confusion: mean confusion score. The final row summarizes these metrics across the entire SATA-Bench dataset.
Data Source n LC m me min max r w FRE FGL ARI DCR GFI Confusion
Reading Comprehension 258 3–15 2.8 2 2 10 na 2018.46 59.94 9.22 12.57 9.27 9.75 0.33
Toxicity 221 8 2.56 2 2 6 4 1015.32 37.83 12.28 13.33 10.49 12.57 0.27
News 248 6 2.36 2 2 5 3 785.93 62.51 8.92 11.15 11.1 10.94 0.26
Laws 281 15 5.3 5 2 10 3 5761.69 45.09 12.29 14.06 8.75 12.07 0.14
Events 202 6 2.63 2 2 5 3 3644.06 50.64 10.83 13.08 9.7 11.8 0.25
SATA-Bench 1470 3-16 3.55 3 2 10 3.2 2491.01 49.56 10.75 12.80 9.96 11.51 0.24
@@ B.1 Initial Filtering
While we cannot entirely eliminate the possibility of memorization, we applied the open-source contamination detection pipeline(Li et al., ). Using the Bing Search API, we found top 20 relevant queries per question to check for verbatim web overlap. We then cross-referenced hits with Common Crawl indexes. We exclude questions that were flagged as contaminated, indicating that our data is neither indexed in Common Crawl nor retrievable via public search. This reduces the likelihood that any model saw our questions during pre-training.
@@ B.2 Human Validation
Human validation is to ensure that the questions are unambiguous. Using humans to validate the question is inspired by  (Tarrant et al., 2006; Moore et al., 2024). For each question in the benchmark, we ask five annotators whether the question contains ambiguous information.
@@
Once it is done, the total cost is tracked (1301.89), with 5 people per label at a cost of 0.012 each. We only select questions that are “Strongly agree” and “Agree" >0.8>0.8.
B.3 Redudency and Consistency Analysis
To ensure the diversity of the dataset labels, we ensure that our answer group has labels with different similarity. To assess label redundancy, we encoded labels using SentenceTransformer (all-MiniLM-L6-v2) and computed pairwise similarities. The mean maximum similarity across label sets is 0.473, with standard deviation 0.206. This confirms a mix of semantically similar and distinct labels. The top 10 percentile score is 0.786 and the bottom 10 percentile score is 0.235. This shows that our dataset has diverse labels with similar percentage of semantically similar and dissimilar labels. Count bias increased after removing similar-label questions, suggesting that LLMs sometimes use semantic similarity to infer related correct answers. We remove all questions that have label pairs with similarity score over 0.786. We then recalculated count bias related metrics across all closed-source models. CtDif is lower and CtDifAbs get higher. This means that removing similar labels in question actually increase the number of count bias. We suspect that is due to the fact that LLM can reasoning through similar labels and use those labels’ similarity to identify all correct answers.
B.4 Human Labeling
To ensure that each question has a valid and correct answer, we conducted a comprehensive human evaluation. An initial manual inspection revealed that some questions lacked clearly correct answers. To verify answer correctness, we recruited three experienced annotators to review all questions that remained after prior filtering and validation. Annotators were compensated at a rate of at least \$35 per hour. Each question was independently evaluated by at least two annotators.
@@ B.4 Human Labeling
Human Labeling Given original answers and LLMs’ answers, you’$l$ try to identify correct answer of the following questions. You’re expected/encouraged to use Google, and any internet resources you can find to try and answer the question correctly. Requirements and Expectations 1. You are encouraged to use Google, and any websites you can think of or find that may help you answer the question and understand the concept. However, you are NOT allowed to use AI assistants like chatGPT, Claude, Grok3 Geimini, etc., or ask people for help. All their answers to the question has been provided anoymously under LLM Answers. 2. We ask that you spend at least 5 minutes trying to answer each question before making your selection. If you haven’t settled on an answer choice in that time, we encourage you to spend as long as you need to be confident in your selection. 3. These questions will be hard, and you will likely need to spend a while on each of them to make progress understanding the context. Read relevant resources, take plenty of time, and answer "I don’t know" if you’re pretty sure you have no realistic way of answering confidently. 4. You will also be given the opportunity to give feedback on the question. We’re especially interested in feedback about whether the question was ambiguous, but please feel free to give feedback of any other form! Suggestions and Strategies for Labeling 1. Look up definitions for all of the unfamiliar terms in the question and answer choices. Keep a list of those definitions handy so you can easily refer back to the definitions if you forget the jargon. 2. LLMs’ answer is not always reliable and original answer is not always correct. Please try to solve the question independently before looking at potential answers. 2. Look for primary resources, like research papers and textbooks, as these can often contain clearer explanations than sources like Wikipedia (although Wikipedia can be useful in many cases as well). You are presented with the following: Paragraph: paragraph Question: question Choices: choice Original Answers: original answer LLM Answers: llm answers Answers:
@@ Appendix C Hyperparameters
GPT‑4‑turbo OpenAI gpt-4o-2024-11-20 11/20/24 OpenAI API
Claude‑3.7 Sonnet Think Anthropic anthropic.claude-3-7-sonnet-thinking-20250219-v1:0 02/24/25 AWS Bedrock
Claude‑3.7 Sonnet Anthropic anthropic.claude-3-7-sonnet-20250219-v1:0 02/24/25 AWS Bedrock
Claude‑3 Sonnet Anthropic anthropic.claude-3-sonnet-20240229-v1:0 02/29/24 AWS Bedrock
@@
Claude‑3.5 Haiku Anthropic anthropic.claude-3-5-haiku-20241022-v1:0 10/22/24 AWS Bedrock
Claude‑3 Haiku Anthropic anthropic.claude-3-haiku-20240307-v1:0 03/07/24 AWS Bedrock
Claude‑3 Opus Anthropic anthropic.claude-3-opus-20240229-v1:0 02/29/24 AWS Bedrock
@@
GPT‑4.1 mini OpenAI gpt-4.1-mini-2025-04-14 04/14/25 OpenAI API
Claude‑3.5 Sonnet Anthropic anthropic.claude-3-5-sonnet-20240620-v1:0 06/20/24 AWS Bedrock
@@
GPT OSS 120B OpenAI openai.gpt-oss-120b-1:0 08/05/25 AWS Bedrock
GPT OSS 20B OpenAI openai.gpt-oss-120b-1:0 08/05/25 AWS Bedrock
@@
Phi‑3 7B Microsoft microsoft/phi-3-small-128k-instruct 05/21/24 Hugging Face
Phi‑4‑mini‑reasoning Microsoft microsoft/phi-4-mini-reasoning 04/15/25 Hugging Face
@@ Appendix D Compute Resources
We use AWS Bedrock batch inference for large models’ inference such as Claude3 Sonnet, Claude 3.5 Haiku, Claude 3 Haiku, Claude 3 Opus, Claude 3.5 Sonnet, Llama 3.1 405B, Mistral Large V2, Llama 3.2 90B, and Llama 3.1 70B. We use AWS cross-region inference for Claude3.7 Reason, Claude3.7, and Deepseek R1. We use official APIs from the respective providers for models such as OpenAI O3, GPT4.1, Grok3 Reason, GPT4, Geimini2.5 Reason, Gemini 2 Flash, GPT 4.1 mini, GPT OSS 120B, GPT OSS 20B, and Qwen Plus.
@@ Appendix E Non-expert Human Benchmark
To contextualise LLM results on SATA‑Bench, we recruited non‑expert annotators on Amazon Mechanical Turk, adapting the instructions from  (Rein et al., 2023). All questions was labelled as follows:
  • Task set‑up. Each question was presented with the original answer options plus decoys (e.g. $\rightarrow\!ABCD$EFGHIJK) to identify inattentive workers. Nine independent annotations were collected per item at a rate of \$0.84 per question, matching the fair‑wage recommendations of GPQA.
  • Quality safeguards. Workers were: (i) informed that every item contains at least two correct answers; (ii) forbidden from consulting LLMs or other people, yet allowed to look up unfamiliar terms on Google/Wikipedia; (iii) required to spend $\geq$2 minutes on each question. Submissions that selected any decoy, took <1<1 min, or violated the lookup policy were discarded (7.1 %).
  • Label selection. From the surviving pool, we randomly drew one annotation as the human label; single‑choice answers were retained to keep the evaluation comparable to LLMs that sometimes return only one option.
@@
Human 17.9 60.6 54.4 45.0 15.3 0.46 1.46 0.6-0.6 34.1 1.44
Table 9: Aggregate performance of crowd annotators on the SATA‑Bench subset.
As anticipated, non‑experts achieve modest exact‑match and precision, yet their selection‑bias metrics (RStd, RSD, SPD) resemble those of mid‑tier LLMs. Crucially, they exhibit smaller absolute count bias ($|\text{CtDif}|$) and higher correct‑count accuracy (CtAcc), indicating superior intuition for the number of correct options even when individual labels are missed. These human baselines therefore offer a realistic point of comparison for evaluating LLM performance on specialised SATA tasks.
@@ F.1 Performance Metrics Definition
  • False Positive Rate (FPR) calculate the fraction of predicted labels that contain labels that are not in the correct labels.
@@ F.2 Selection Bias Metrics Definition
    where $k$ is the number of choices, $r_{i}$ is the recall of the $i$-th class, and $\bar{r}$ is the arithmetic mean of $r_{i}$ values. Note that our recalls are calculated at the label level since this is multi-class question (Zheng et al., 2024)
@@
    where $k$ is the number of choices, $s_{i}$ is the accuracy of the $i$-th class, and $\bar{s}$ is the mean accuracy averaged across classes. Please note that our recalls are calculated at the label level since this is multi-class questions (Croce et al., 2020; Reif & Schwartz, 2024)
@@ Appendix G Unselection bias metric
SPD has a minimal value of 0 at $q_{i}=p_{i}$ for all $i$, when the prediction aligns with the ground truth. SPD diverges as $q_{i}\rightarrow 0$ while $p_{i}$ is finite for any $i$, when the LLM shows a non-selection bias against a particular label. SPD also diverges as $p_{i}\rightarrow 0$ while $q_{i}$ is finite for any $i$, when the LLM shows a selection bias toward a particular label. In this sense, SPD serves as a metric to measure the disagreement of choice probability between the ground truth and the prediction, reflecting both under-selection and over-selection. (See Appendix G.2 for the mathematical analysis.)
@@ G.1 Behavior of SPD Metric
We repeat the above sampling process $M$ times, and compute the True rate of each option for the ground truth $p_{i}$} and the prediction $q_{i}$}, with $i=1,2,3,4$. We then substitute the numbers into Eq. equation 3 to calculate SPD. Note that in the current setting, $p_{i}=\mathtt{p},\forall i$, and $q_{1}=\mathtt{q}$, $q_{2,3,4}=\mathtt{p}$.
@@
Figure 8: Relationship between Selection Probability Divergence (SPD) and prediction probability ($\mathtt{q}$) across different ground truth probabilities ($\mathtt{p}$). The curves are averaged over 100 replicates, and the shaded area represents the standard deviation. In each plot, the minimal value of SPD is 0 at $\mathtt{q}=\mathtt{p}$, when the prediction aligns with the ground truth.
@@ G.2 Sensitivity of SPD to label probability ratio
We analyze the behavior of SPD as the relationship between $p_{i}$} and $q_{i}$} changes. We first define the ratio of the two probabilities as $r_{i}\equiv q_{i}/p_{i}$, $i=1,2,\ldots,k$, and rewrite the SPD definition Eq. equation 3 as
@@
As the misalignment between the ground truth and the prediction grows, either with $r_{i}\rightarrow 0$ or $r_{i}\rightarrow+\infty$, SPD diverges according to Eq. equation 4. Therefore, a large value of SPD reflects the disagreement of the choice probability between the ground truth and the prediction.
@@ H.2.1 Problems for current MCQ prompts
Existing benchmarks (Intelligence, 2024) use the following prompts for MCQ questions and then use exact match to get the correct option.
@@
We can observe that parsing the above prompts is much harder because we do not know when they will output the option ID plus answers, and we cannot blindly extract all matched option IDs. Following (Intelligence, 2024) page 38, we used the JSON format to extract answers. Closed-source LLMs have shown good instruction following capability for JSON format output (Zhou et al., 2023). Also, python has many packages that can do fuzzy JSON matching, which can be used to handle edge cases. We extract as many JSON files as possible from each output and combine the extracted answers together as the final answer to the question.
SATA Prompt Given the following question where there is more than one correct answer, choose all correct answers. Question: question Choices: choices Please select all choices that apply. You must focus on the question and select all choices that apply. Let’s think step by step: You must present your selected option IDs in the following JSON format: $\{{"choices":<A|B|C|D|E|F|G|H|I|J|K|L|M|N|O>}\}$
SATA Prompt Example Given the following question where there is more than one correct answer, choose all correct answers. Question: What is the correct answer to this question: Bangladesh passed on its weekend tender for 100,000 tones of optional origin soft wheat, private export sources said. What topics are related to the document above?Choices: A. Wheat, B. Barley, C. Grain, D. Retail, E. Money-supply, F. Sugar. Please select all choices that apply. You must focus on the question and select all choices that apply. Let’s think step by step: You must present your selected option IDs in the following JSON format: $\{{"choices":<A|B|C|D|E|F|G|H|I|J|K|L|M|N|O>}\}$ Answer (from GPT4o): Let’s analyze the question step by step: Understanding the question context: The document discusses Bangladesh’s weekend tender for 100,000 tonnes of "optional origin soft wheat." Analyzing the options provided: A. Wheat: The document specifically mentions "soft wheat." This choice is correct. B. Barley: There is no mention of barley in the document. This choice is incorrect. C. Grain: Wheat is a type of grain. This choice is correct as the broader category includes wheat. D. Retail: The document does not mention retail activities. This choice is incorrect. E. Money-supply: There is no discussion of money supply or related economic terms in the document. This choice is incorrect. F. Sugar: The document does not mention sugar. This choice is incorrect. Conclusion: Based on the analysis, the related topics are Wheat (A) and Grain (C). Final Answer in JSON format: { "choices": ["A", "C"] }